OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the point (−1, 2) and has a slope of 0.9. y = x + i think m=0.9 and y=2

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

how did u get y intercept = 2 ?

OpenStudy (anonymous):

i thought you had to use the intercept in the point given

ganeshie8 (ganeshie8):

Nope. A point has no intercept. We define intercepts for graphs like lines, parabolas etc..

ganeshie8 (ganeshie8):

\[\large \rm y = 0.9x + b\]

ganeshie8 (ganeshie8):

plugin the given point (-1, 2) in above equation and find out the y intercept, \(\rm b\)

ganeshie8 (ganeshie8):

plugin x = -1 and y = 2

OpenStudy (anonymous):

y=0.9x+-1 y=0.9x+2

OpenStudy (anonymous):

i did that wrong

ganeshie8 (ganeshie8):

Not quite, you need to plugin x = -1 and y = 2 in the equaiton and find out the value of \(\rm b\) first. Let me show you how to work it

OpenStudy (anonymous):

you get 1.8 and -0.9

ganeshie8 (ganeshie8):

You have this equation :\[\large \rm y = 0.9x + b\] plugin \(\rm x = -1\) and \(\rm y = 2\), you get : \[\large \rm 2 = 0.9(-1) + b\]

ganeshie8 (ganeshie8):

\[\large \rm 2 = -0.9 + b\]

ganeshie8 (ganeshie8):

\[\large \rm 2 +0.9 = b\]

ganeshie8 (ganeshie8):

\[\large \rm 2 .9 = b\]

ganeshie8 (ganeshie8):

So the y intercept is \(\large \rm 2.9\)

OpenStudy (anonymous):

oh, okay. that makes sense.